OpenStudy (anonymous):

Let y = 32,000(0.94)x represent the purchasing power of $32,000 with an inflation rate of 6% per year, where x represents the number of years. What is the purchasing power when x = 0? Answer $94,000 $30,080 $32,000 $64,000

OpenStudy (anonymous):

Do I simplify first orr?

OpenStudy (anonymous):

@phi cam you help me?

OpenStudy (anonymous):

Thanks for helping phi..

OpenStudy (phi):

you probably mean \[ y = 32,000(0.94)^x \] you should write this with an up arrow to show x is an exponent y = 32,000(0.94)^x they want you to replace x with a 0 (that means change x to 0 everywhere you see it in the equation) : \[ y = 32,000(0.94)^0 \] now simplify using the important fact that any number to the 0 power is 1

OpenStudy (anonymous):

okay give me a sec

OpenStudy (anonymous):

Would the answer be 32000?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@zordoloom is my answer correct sir?

OpenStudy (anonymous):

@phi sir I got 32000 for my final answer..

OpenStudy (anonymous):

Yes that seems correct.

OpenStudy (phi):

yes, it is 32000* (.94)^0 but 0.94^0 is 1 so this is just 32000*1= 32000

OpenStudy (anonymous):

Thanks guys! (:

OpenStudy (phi):

Common sense should tell you that inflation after 0 years (right now!) should have no effect... it takes time for the money to lose its value